Subject 2: Oligopolistic Competition (35% Two firms (Natural Salt and Healthy Salt) compete i Consumers see the salt produced by both firms as perfect substitutes. n the market for Himalayan table salt In this market, each firm chooses what output to produce and price is determined by aggregate output. Market demand is given by Q 450-2P, where Q is kgs ard Pis €/kg. The initial marginal cost of Natural Salt is S0 kg. The respective for Healthy Salt is 40/kg. A process innovation in the proxduction technology of Himalayan table salt would reduce the Natural Salts marginal cost to 44 and Healthy Salts marginal cost to 34 (a) Consider that only one of the two firms can introduce the aforementioned process innovation in its production technology, while its rival firm still produces with its initial marginal cost. How much would each firm be willing to pay so as to introduce the aforementioned process innovation in its production technology? IMark 1.0 (b) Consider that each firm has the option to pay 1,000 euros so as to introduce the aforementioned process innovation in its production technology. Given each firms decision, on whether to introduce the innovation or not, firms compete in the market by choosing simultancously their level of output. If you were the Natural Salts manager would you suggest the introduction of the innovation or not? What would you suggest if you were the Healthy Salts manager? (Hint: Construct a payof matrix.] IMark 1.5 (c) Assume that Natural Salt and Healthy Salt decide on the following: Healthy Salt pays 1.000 euros and introduces the process innovation. Then, the two firms merge and produce with marginal cost equal to 34. Healthy Salt keeps 500 euros of the post-merger firms ess innovations cost) and the remaining profits are equally shared profits (half of the proc to the merger participating firms. Do both firms have incentives for such a merger compared to the equilibrium in (b)? [Mark 0.5 (d) How does the merger change consumers surplus and social welfare, compared to t respective ones in (b)? IMark 0.5 [Marking scheme: economic intuition 10%, use of appropriate form ulae 40%, correct calculations 30%, overall presentation 20%)
Answer #1

良= 450-1? anut diniane MCH = 34 2 6, (yp-o.th) _ 44。 2 2 2 2 2 ラ t 202 = 370 2. 70 I 2

ー一45で 2 o, + 202-382 LA9 2 ( 382-20) + O2 = 350 to 4 x 118 6a 62 5 (13 (103メ126)-(Igix 4の (103 x 198)÷ (gux133)/ S 2238 1 53 y

206 2 (382-202) +. OL-362 Natuural alt Heeuhs Health낭 o·S· か^.tqRLか..J. WP.

ニ 95.5 ニ 15, 128 2 一 58 79-78 IT as ann

DWL 282 x (9ss)-34 ) x (382 ~ 259)二3782,25 in Cb) . DwLu 二.lx(101-94)x. ( 3 62-lly 2